OpenStudy (anonymous):

Solve the equation for y: 4x + 3y = -12.

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

y = -12 - 4/3 x y = -4 - 4/3 x y = -12 + 4x y = -4 + 4x

OpenStudy (anonymous):

is it the 3rd one?

OpenStudy (skullpatrol):

No.

OpenStudy (anonymous):

uhh the last one

OpenStudy (e.cociuba):

Well first why don't you stop guessing and let @skullpatrol walk you through it... And second start solving the problem. Start by telling us what you have to do first and we will help you along the way. :)

OpenStudy (e.cociuba):

Ok so what do you know? Do you have any idea how to start?

OpenStudy (skullpatrol):

You can take this one @e.cociuba thanks :D

OpenStudy (e.cociuba):

Np.

OpenStudy (e.cociuba):

Ok so @rasta420 What do you think you have to do first?

OpenStudy (e.cociuba):

Just looking for the answer? Well thats not how I role. Have a great day! :)